Let L be the line with equation y=x-1. Find the minimum distance between L and the point (4,5) by using (a) theorem 10.2.4, (b) the Method of Lagrange.
1
Expert's answer
2021-07-18T11:04:55-0400
"d^2=(x-4)^2+(y-5)^2"

(a) Given "y=x-1."

Then


"d^2=f(x)=(x-4)^2+(x-1-5)^2"

"f'(x)=2(x-4)+2(x-6)"

"=2(x-4+x-6)=4(x-5)"

Find the critical number(s):


"f'(x)=0=>4(x-5)=0=>x=5"

Critical number: "5."


"f''(x)=4>0"

Then the function "f(x)" has a local minimum at "x=4."

Since the function "f" has the only extremum, then the function "f(x)" has an absolute minimum at "x=4."


"x=5, y=5-1=4"

"d=\sqrt{(5-4)^2+(4-5)^2}=\sqrt{2}"

(b) Given "y=x-1."

Then


"L(x,y)=(x-4)^2+(y-5)^2-\lambda(x-y-1)"

"\dfrac{\partial L}{\partial x}=2(x-4)-\lambda"

"\dfrac{\partial L}{\partial y}=2(y-5)+\lambda"

"\dfrac{\partial L}{\partial \lambda}=-x+y+1"


"\dfrac{\partial L}{\partial x}=0"

"\dfrac{\partial L}{\partial y}=0"

"\dfrac{\partial L}{\partial \lambda}=0"




"\lambda=2(x-4)"

"y=x-1"

"2(x-1-5)+2(x-4)=0"


"x=5"

"y=4"

"\lambda=2"

The minimum distance between L and the point (4,5) is "\sqrt{2}."